### v2

# Evaluation and Facilitation of Online Discussions in the LLM Era: A Survey

Abstract
We present a survey of methods for assessing and enhancing the quality of online discussions, focusing on the potential of Large Language Models (LLMs). While online discourses aim, at least in theory, to foster mutual understanding, they often devolve into harmful exchanges, such as hate speech, threatening social cohesion and democratic values. Recent advancements in LLMs enable artificial facilitation agents to not only moderate content, but also actively improve the quality of interactions. Our survey synthesizes ideas from Natural Language Processing (NLP) and Social Sciences to provide (a) a new taxonomy on discussion quality evaluation, (b) an overview of intervention and facilitation strategies, (c) along with a new taxonomy of conversation facilitation datasets, (d) an Large Language Model (LLM)-oriented roadmap of good practices and future research directions, from technological and societal perspectives.
Evaluation and Facilitation of Online Discussions in the LLM Era:
A Survey
[…]
1 Introduction
Discussions, especially of complex or controversial topics, are a cornerstone of collective decision-making Burton et al. (2024). In contrast to initial hopes of promoting mutual understanding Rheingold (2000), online discussions (especially in social media) often degenerate into hate speech, personal attacks, promoting conspiracy theories or propaganda – to the extent that they can even be considered a threat to social cohesion and democracy Tucker et al. (2018); Mathew et al. (2019).
Natural Language Processing (NLP) and Machine Learning (ML) can potentially help improve the quality of online discussions. For example, automatic classifiers Bang et al. (2023); Molina and Sundar (2022) are already being used to help or even replace human moderators, by flagging posts that violate the law or policies of online discussion fora Saeidi et al. (2021).
Social Science provides theories and applications for the facilitation of a discussion, but in specific contexts, such as teaching/learning Mansour (2024) or clinical discussions Gelula (1997), without much research devoted to online discussions, such as in social media. While prior NLP studies have explored LLM-facilitated discussions Burton et al. (2024); Aher et al. (2023); Beck et al. (2024); Schroeder et al. (2024); Small et al. (2023); Cho et al. (2024), rarely does Social Science work examine how facilitation can be automated Gimpel et al. (2024).
Figure 1: A conceptualization of this survey. We explore approaches from different disciplines, which recommend their own ways of evaluating and improving discussions.
In this survey, we combine LLM-based methods, with ideas from Social Science (e.g., Deliberative Theory) when discussing how to evaluate online discussions, and when exploring intervention strategies. Figure 1 provides a high-level conceptualization of our work.
The main research question of this survey is can LLMs be used effectively as facilitators in online discussions? To explore this question, we focus on three key areas: (1) methods (potentially also LLM-based) for evaluating aspects of online discussions, (2) intervention strategies for facilitation, and (3) available data resources relevant to facilitation. Specifically, we survey discussion evaluation aspects and introduce a new taxonomy (§4). We map tasks suited for ML models, LLMs, and humans, aggregate multidimensional insights on facilitation strategies (§5), and outline future possibilities for LLMs6). Additionally, we compare major datasets, dividing them into categories per task (§7). Our work focuses mostly on written thread-like discussions (§2).
Our findings show that (a) many discussion evaluation dimensions coexist in the literature; (b) LLM advancements show significant promise in improving the quality and timeliness of facilitation methods; (c) while surveying the existing datasets, we notice a scarcity of datasets for studying facilitation. We posit that LLM-generated discussions, could become an asset to develop and test automatic facilitation strategies in diverse artificial discussions, before testing the strategies and the LLM-based facilitator agents in more costly experiments with human participants.
2 Terminology
Given the numerous aspects to consider regarding discussion quality and facilitation, we clarify the terminology we use. We highly recommend consulting the Terminology Section of Appendix C and, especially, Table 3, where we explain our findings with regard to the terms used in the literature.
Facilitation vs. Moderation
The term ‘moderation’ is more commonly used in NLP Argyle et al. (2023), typically referring to the flagging and/or removal of unwanted content (‘content moderation’), while ‘facilitation’ is more prevalent in the Social Sciences, where it encompasses a broader scope, including active interventions Vecchi et al. (2021); Kaner et al. (2007); Trenel (2009). Given the limited attention to facilitation in NLP and the survey’s grounding in Social Science, we distinguish between the terms, even though they are sometimes used interchangeably in the literature.
Ex-Post moderation
This survey mainly focuses on ‘Real-Time, Ex-Post-moderation’, i.e., moderation happening just after the user has posted some content. This is different from pre-moderation approaches, such as nudging users before they post harmful content Argyle et al. (2023), or delaying the posting of user content until a moderator has had the chance to check it.
Discussion, Deliberation, Dialogue, Debate
The definitions of these terms often vary across literature Russmann and Lane (2016); Goñi (2024). We focus on discussions, a general term for verbal/written exchanges Russmann and Lane (2016), and deliberations, a term for structured discussions focusing on opinion sharing Degeling et al. (2015); Lo and McAvoy (2023). This is in contrast to the (at least in theory) collaborative nature of dialogues Rose-Redwood et al. (2018); Bawden (2021); Goñi (2024) and the competitive and organized nature of debates Lo and McAvoy (2023).
Tree-style discussions
(or “threads”) are discussions which start from an Original Post (OP) with subsequent comments replying to either the OP or to other comments Seering (2020).
3 Comparison to Other Surveys
Only two studies have surveyed the field of NLP while also considering ideas from Social Science. However, they focus mainly on Argument Mining (AM). These are the studies of Wachsmuth et al. (2024) and Vecchi et al. (2021). Wachsmuth et al. (2024) focus primarily on discussion evaluation disregarding its relation to facilitation, which is one of the main goals of our survey. The survey of Vecchi et al. (2021) argues that advancing AM for social good requires a collaborative effort between AM and Social Science. They point out that traditional AM has prioritized the logical structure and soundness of arguments, while overlooking other important dimensions, such as civility, respectfulness, inclusiveness, originality, and the broader impacts of discussions—such as encouraging mutual understanding and problem-solving. Building on these notions, we incorporate ideas from Social Science into NLP-based approaches, discussing both discussion evaluation and facilitation, both with a focus on the potential of LLMs.
4 Discussion Quality Evaluation
Improving online discussions presupposes being able to define and measure discussion quality. While there have been attempts to provide frameworks for discussion quality evaluation Kies (2022); Gerber et al. (2018), none of them is directed towards facilitation. Crucially, most existing frameworks ultimately rely on human judgments as their reference point, yet human evaluation is expensive, slow, and shows low inter-rater agreement on dimensions that involve subjective interpretation, such as pragmatic cues (Smith et al., 2022; Yeh et al., 2021; Khalid and Lee, 2022). This evaluation bottleneck motivates a taxonomy of evaluation methods that is both comprehensive and amenable to scalable automatic measurement.
In this work, we draw from the works of Bächtiger et al. (2022, 2010); Steenbergen et al. (2003); Falk and Lapesa (2023) and Kies (2022) to define a new social-science-informed taxonomy for discussion quality dimensions. While we present a structured taxonomy, it is important to note that the categories are not mutually exclusive. Rather, elements within the taxonomy may coexist within evaluation dimensions, complement one another, or serve as explanatory mechanisms for other dimensions. An example of the dimension interaction can be found in Table F in the Appendix. The grouped dimensions along with the NLP approaches are shown in the Appendix in Table 4.
4.1 Structure and Logic
Argument Structure and Analysis
Argument Quality (AQ) is a multidimensional concept assessed through logical, rhetorical, and dialectical dimensions Wachsmuth et al. (2017). The logical dimension focuses on the coherence and structure of the argument. The rhetorical dimension assesses persuasiveness, focusing on the argument’s style and emotional appeal. The dialectical dimension assesses the constructiveness of the argument. Empirically, threads with well-formed claim-evidence chains exhibit higher coherence and lower odds of devolving into ad-hominem attacks, making AQ scores, as a discussion quality dimension, an early-warning indicator of derailment (Chang and Danescu-Niculescu-Mizil, 2019). All the above dimensions of automatic argument-structure analysis can be used by a facilitator to keep the discussion fact-centered, inclusive, and on track Falk et al. (2021); Falk and Lapesa (2023).
Coherence and Flow
‘Coherence’, as described above, evaluates logical consistency, while ‘flow’ assesses smooth progression in discussions Li et al. (2021). Both are essential tools for facilitators in their effort to redirect off-topic comments and guide transitions between topics during a discussion Lambert et al. (2024); Park et al. (2012); Falk et al. (2024). A sudden drop in how well responses match the topic or question often comes before personal attacks or off-topic turns (Chang and Danescu-Niculescu-Mizil, 2019; Zhang et al., 2018), making coherence and flow indicators of argument structure and a valuable early signal for facilitators.
Turn-taking
How speakers alternate, the frequency of their turns, and the participants they address can serve as a diagnostic of conversational health. Balanced exchanges enhance coherence (Cervone and Riccardi, 2020), predict constructiveness (§4.3) (Niculae and Danescu-Niculescu-Mizil, 2016), and provide facilitators with actionable cues (Schroeder et al., 2024). To gauge speaking time, turn count, and word usage, researchers have applied metrics such as entropy Niculae and Danescu-Niculescu-Mizil (2016) and Gini coefficients Schroeder et al. (2024).
Linguistic Markers
Linguistic markers have been used to help model content and expression in online discussions Wilson et al. (1984). Early methods used lexicons for sentiment, toxicity, politeness (§4.2 and 4.3) and collaboration evaluation Lawrence et al. (2017); Avalle et al. (2024). For example, spikes in hedges (e.g., ‘maybe’, ‘I guess’) invite clarification requests by facilitators, while bursts of second-person pronouns, similarly to turn-taking, often foreshadow personal attacks and can prompt a civility nudge Niculae and Danescu-Niculescu-Mizil (2016).
Speech and Dialogue Acts
Rooted in Speech Act Theory Austin (1975); Searle (1969), dialogue acts have been employed to assess deliberative quality and analyze facilitation strategies Fournier-Tombs and MacKenzie (2021); Chen et al. (2024a). They characterize dialogue turns (e.g., interruption) to analyze interaction dynamics Ferschke et al. (2012); Stolcke et al. (2000); Zhang et al. (2017); Al-Khatib et al. (2018). Positive (e.g., causal reasoning) or negative (e.g., disrespect) dialogue acts can be scored to reflect discussion quality and low scores may indicate the need for interventions Ziems et al. (2024); Cimino et al. (2024); Martinenghi et al. (2024); Schroeder et al. (2024).
Pragmatic Comprehension
Pragmatic comprehension—how context shapes meaning—is crucial to facilitation, as intended meanings often diverge from literal expressions (i.e., implicature). Humans resolve such ambiguity using social and commonsense knowledge. Grice’s maxims Grice (1975), a central pragmatic concept, can help explain this process by outlining the conversational principles people rely on to infer meaning, while they have already been used to assess discussion quality Jwalapuram (2017); Langevin et al. (2021); Ngai et al. (2021); Nam et al. (2023).
4.2 Social Dynamics
Politeness
Politeness serves as a cornerstone of prosocial behavior, an attribute that facilitators desire to foster in online discussion forums Lambert et al. (2024). In the context of facilitation, it has mainly been studied in relation to conversational derailment (§7) Zhang et al. (2018) and constructiveness (§4.3) De Kock and Vlachos (2021); Zhou et al. (2024).
Power and Status
Power and status influence conversational dynamics, affecting language use and turn-taking (§4.1). Higher status speakers can control the flow of discussions and foster social inequalities. Interestingly, low-status individuals tend to mimic the linguistic styles of high-status speakers more than the opposite Danescu-Niculescu-Mizil et al. (2012), and this can be used as a signal that there is high/low-status imparity in a discussion. Facilitators may intervene, then, to ensure that the right to speak is evenly distributed among participants, preventing projection of social biases and stereotypes.
Disagreement
Disagreements, when constructive, improve discussions by fostering deeper understanding Friess (2018); De Kock and Vlachos (2021). Assessing disagreement, however, is complex. The hierarchy of Graham, 2008 considers disagreement tactics ranging from name calling to refuting the central point. Along with other work on dispute tactics Walker et al. (2012); Benesch et al. (2016); De Kock et al. (2022), it can be used to examine types of disagreements in a discussion.
4.3 Emotion and Behavior
Empathy
Empathy is the ability to understand others’ perspectives and emotions and respond correspondingly Lipman (2003); Xu and Jiang (2024). Facilitators desire to foster empathy in online discussions, since it encourages prosocial behavior and boosts engagement Xu and Jiang (2024); Concannon and Tomalin (2024); Lambert et al. (2024). To do so, they encourage users to share personal stories and experiences Schroeder et al. (2024). Various coding schemes Macagno et al. (2022), psychological indicators (e.g., the emotion-laden words of Furniturewala and Jaidka, 2024), and dimensions (e.g., perceived engagement such in Xu and Jiang, 2024) have been used to detect both expressed and perceived empathetic traits.
Toxicity
Toxicity in online discussions refers to harmful or disrespectful language that hinders productive discourse and can derail meaningful discussions Avalle et al. (2024). Facilitation is key to maintaining healthy communication, requiring both early detection of toxicity and (in the case of more active facilitation) proactive de-escalation strategies, such as conversation redirection or positive engagement (§5). In the case of conventional moderation that only aims to flag or remove toxic content, debate persists over what content warrants removal Warner et al. (2025); Habibi et al. (2024); Pradel et al. (2024).
Sentiment
Sentiment analysis helps identify whether discussions are positive, negative, or neutral. In the context of facilitation, sentiment analysis gauges the tone of discussions, which influences the quality of interactions De Kock and Vlachos (2021). Positive sentiment contributions in online discussion forums usually signal prosocial behavior and hence are highly encouraged by facilitators Lambert et al. (2024), while negative sentiments among discussants contribute to conversation toxicity Avalle et al. (2024).
Controversy
Controversy arises from divergent viewpoints, leading to polarized exchanges that can escalate to toxicity and derail online discussions Avalle et al. (2024). Controversial comments have been shown to contribute to a decline in positive emotions and a sustained rise in anger Hessel and Lee (2019); Chen et al. (2024b). The spread of political leanings among discussants and sentiment distribution analysis are common approaches to measure controversy Avalle et al. (2024).
Constructiveness
Constructiveness fosters meaningful dialogue, especially in online discussions, by promoting resolution and cooperation Shahid et al. (2024). It is often signalled by linguistic markers (§4.1) De Kock et al. (2022); Falk et al. (2024). A facilitator can exploit a constructiveness score; threads trending upward are worth highlighting or summarizing, whereas a downward drift may trigger facilitation tactics such as slower, structured turn-taking or clarification prompts (De Kock and Vlachos, 2021).
4.4 Engagement and Impact
Engagement
Engagement is desirable in online discussion platforms as it combines interest and participation Lambert et al. (2024); Park et al. (2012). It is proxied by measures like reciprocity Graham and Witschge (2003); Stromer-Galley (2007); Zhang et al. (2018), number of comments posted by each user Avalle et al. (2024), discussion length Adomavicius (2021); Avalle et al. (2024), while Ferron et al. (2023) define subdimensions such as response diversity, interestingness, and specificity.
Persuasion
Empirical literature has primarily examined factors influencing persuasion that align with other categories in our taxonomy, such as linguistic markers (§4.1) and turn-taking (§4.2)Tan et al. (2016). Considering this connection, persuasion is not only an indicator of argument quality, but may also serve as a proxy for identifying additional markers signaling whether facilitator intervention is needed.
Diversity and Informativeness
Diversity in online discussions refers to the presence of varied perspectives, backgrounds, and experiences, which can enrich conversations by fostering constructive exchanges Irani et al. (2024); Zhang et al. (2024). To prevent echo chambers and promote inclusivity, facilitators can use diversity measures to encourage opinion diversity Anastasiou et al. (2023), encouraging users to explore a broad range of perspectives on a given issue Kim et al. (2021). Informativeness refers to the relevance and value of information shared in a discussion and is considered a building stock of prosociality, an attribute that facilitation trys to foster in online discussion platforms Lambert et al. (2024).
4.5 LLM Approaches to Discussion Quality
LLMs can significantly aid in evaluating discussion quality, performing on par with humans in annotating argument structure, coherence, and flow across tasks like argument mining and synthesis Mirzakhmedova et al. (2024); Rescala et al. (2024); Wang et al. (2023); Chen et al. (2024a); Irani et al. (2024); Anastasiou and De Liddo (2024); Zhang et al. (2024); Mendonca et al. (2024); Zhang et al. (2023). LLMs can also reliably label dialogue acts Ziems et al. (2024); Cimino et al. (2024); Martinenghi et al. (2024); Schroeder et al. (2024), as well politeness, power, disagreement, and toxicity Zhou et al. (2024); Ziems et al. (2024). However, they struggle with tasks involving social norms (such as irony and humor) and show limited accuracy in sentiment and engagement detection Hu et al. (2023); Sravanthi et al. (2024); Furniturewala and Jaidka (2024); Xu and Jiang (2024). While LLMs show promise in measuring controversy and persuasion, performance drops at the discussion level, especially when aiming to measure diversity, informativeness, and generally when social and pragmatic understanding is necessary Ziems et al. (2024); Avalle et al. (2024); Lawrence and Reed (2020).
5 Intervention Strategies
5.1 When to Intervene
Picking the right moment to intervene is a crucial part of effective facilitation strategies. If a facilitator does not intervene when they should have, there is a risk of significant escalation, while intervening when unnecessary can increase toxicity Schaffner et al. (2024); Trujillo and Cresci (2022); Schluger et al. (2022); Cresci et al. (2022). It is imperative then (also considering the evaluation dimensions discussed in §4), for a facilitator to be able to recognize subtle cues hinting towards escalation, in order to defuse the situation, something that even experienced human facilitators are not confident to reliably do Schluger et al. (2022). The NLP task of ‘Conversational Forecasting’ may contribute towards this direction. Given a conversation up to a point, a model attempts to predict if an event will occur in the future in that conversation. In our case, this is where a facilitator would intervene Schluger et al. (2022). Traditional ML models can perform well on this task, although their performance varies Falk et al. (2021); Park et al. (2012); Falk et al. (2024); Schluger et al. (2022).
5.2 How to Intervene
There is currently no agreed-upon taxonomy for facilitator interventions. Lim et al. (2011) propose a taxonomy that focuses on discussion facilitation, excluding, however, disciplinary or administrative actions, which are common in online discussions. Park et al. (2012) propose another taxonomy consisting of seven moderator functions, ranging from policing the discussion to solving technical issues. These functions roughly correlate with the volunteer moderator roles, as described by Seering (2020). More practical approaches can be found in facilitator manuals eRulemaking Initiative (2017); MIT Center for Constructive Communication (2024) and books White et al. (2024).
Facilitators often have to decide what form of coercive measure to take to make sure the conversation remains healthy, without having to intervene repeatedly. Human interventions typically use an unofficial ‘escalation ladder’ (Figure 1), where the facilitator will progressively move from milder facilitation tactics to threatening, and finally disciplinary action Seering (2020). ‘Conversational moderation’ Cho et al. (2024), where a facilitator first converses with the offender, has proven effective and is actively encouraged in some facilitator guidelines The Commons (2025). This is probably why disciplinary action is typically not the first choice of a facilitator Schluger et al. (2022) and why it should reasonably be used as a last resort. Softer kinds of interventions that facilitators frequently use first include: setting and informing users about rules Schluger et al. (2022); Seering (2020), welcoming new users Schluger et al. (2022), summarizing key points Small et al. (2023); Falk et al. (2024), balancing participation Kim et al. (2021); Fishkin et al. (2018), and aiding users improve their points Tsai et al. (2024); Falk et al. (2024).
5.3 Personalized Interventions
It is worth stressing that intervention strategies should not be applied en masse, without considering the characteristics of each individual. Traditionally, massive application (or threatening) of disciplinary action has led to adverse effects community- and platform-wide Trujillo and Cresci (2022); Falk et al. (2021) and the creation of echo-chambers Cho et al. (2024). There are also calls for research to move away from one-size-fits-all approaches and instead move towards personalized interventions Cresci et al. (2022). Human facilitators are often able to personalize interventions per individual Schluger et al. (2022), and we hypothesize that LLMs can also do so to some extent.
6 Towards LLM-based facilitation
Until recently, ML models used as facilitation agents were confined to either performing menial tasks, such as pasting automated messages Seering (2020); Schluger et al. (2022), suggesting facilitation actions (e.g., rejecting posts), possibly via human-in-the-loop frameworks Fishkin et al. (2018); Gelauff et al. (2023), identifying possibly escalatory comments Schluger et al. (2022), or employing pre-programmed facilitative tactics, as in the work of Kim et al. (2021), where the model produces automated messages encouraging participation. However, older ML-based and rule-based facilitation are not effective enough to meet the high demands of most platforms Seering (2020); Schaffner et al. (2024).
Advances in LLMs enable the development of facilitation agents that more actively engage in discussions. These agents can warn users about policy violations Kumar et al. (2024), suggest rephrasings to improve tone or persuasiveness Bose et al. (2023), monitor turn-taking Schroeder et al. (2024), and summarize or visualize key discussion points Small et al. (2023). They can also assist in drafting group statements that reflect diverse viewpoints Tessler et al. (2024). A brief, non-exhaustive summary of the capabilities of simpler ML models, LLMs, and humans can be found in Figure 2.
Figure 2: Capabilities of simpler ML, LLM, and human facilitation. Task complexity and cost increase from left to right. Intermediate tasks are handled suboptimally by the preceding method.
6.1 Administrating the Discussion
LLMs are able to tackle a variety of ‘administrative’ facilitation tasks that help structure discussions. For example, facilitators often summarize the views of the participants, seek confirmation of understanding, and share perspectives. This iterative summarization is a task LLMs may handle effectively Small et al. (2023); Burton et al. (2024). However, Feng and Qin (2022) point out some challenges such as discussions with multiple participants, topic drifts, multiple co-references, diverse interactive signals, and diverse domain terminologies. Still, according to Jin et al. (2024), LLMs bring significant advantages over conventional ML methods, “notably in the quality and flexibility of the generated texts and the prompting paradigm to alleviate the cost of training deep models”.
In some deliberative contexts, facilitators are also encouraged to begin a discussion with their own opinion Small et al. (2023), although others disagree MIT Center for Constructive Communication (2024). This is a task LLMs can also handle, albeit less convincingly than Information Retrieval (IR) approaches Karadzhov et al. (2021).
Finally, LLMs can help marginalized groups in discussions by offering translations of the discussions in their native languages, and by helping them phrase their opinions with proper grammar and syntax Tsai et al. (2024); Burton et al. (2024). This can directly improve discussions by increasing their diversity (Section 4.4).
6.2 Evolving Traditional Automation Models
LLMs have been proven to be adept at NLP tasks such as the detection of hate speech Shi et al. (2024), toxicity Kang and Qian (2024); Wang and Chang (2022), and misinformation Kang and Qian (2024); Wang and Chang (2022). These abilities make LLMs usable as drop-in replacements for traditional ML models for these tasks, suggesting that conversational LLM facilitation agents may be able to identify, and dynamically adapt to such phenomena properly. We note however that LLMs are much more expensive and less scalable than their simpler ML counterparts. Furthermore, LLM annotation has its own challenges: LLM survey responses Jansen et al. (2023); Bisbee et al. (2024); Neumann et al. (2025) and annotations Gligori’c et al. (2024) are generally unreliable and surface-level. Non-deterministic behavior is also common in LLMs Atil et al. (2025), but also particularly in closed-source models Bisbee et al. (2024) on which a lot of research on LLM annotation hinges.
6.3 Fully Automated LLM-based Facilitation
There are indications that LLMs can be used as facilitators in the fullest capacity of the role. LLMs are able to predict optimal facilitation tactics Schroeder et al. (2024), like traditional ML models Al-Khatib et al. (2018). Furthermore, they have proven capable of developing and executing social strategies in other tasks, e.g., negotation games, LLM interactions Abdelnabi et al. (2024); Cheng et al. (2024a); Martinenghi et al. (2024). Given that relatively simple ML chatbots, which do not leverage generative text capabilities, have been reported to improve discussions Kim et al. (2021), many expect LLM-based facilitation to be a promising solution to the well-known bottleneck of human facilitation Small et al. (2023); Seering (2020); Burton et al. (2024); Schroeder et al. (2024). Notably, Cho et al. (2024) successfully use LLM facilitators with prompts based on Cognitive Behavioral Therapy to moderate a live discussion with human participants. Their work shows that LLM facilitators can adapt in their instructions to users, although they cannot by themselves affect the discussion with regard to cooperation and mutual respect between the participants.
Nevertheless, LLMs have inherent limitations that make them worse than humans in most social tasks (Figure 2) Rossi et al. (2024). While human facilitators are encouraged to be neutral White et al. (2024); eRulemaking Initiative (2017), numerous studies point to biases in sociodemographic, statistical, and political terms in LLMs R.Anthis et al. (2025); Hewitt et al. (2024); Rossi et al. (2024), which can be exacerbated during the course of a discussion Taubenfeld et al. (2024).
7 Facilitation Datasets
In this section, we provide an overview of the most prominent datasets for online facilitation, considering their sizes and their relevance to core facilitation tasks. We propose the following new taxonomy of facilitation datasets: Conversation Derailment datasets, where the task is to predict when a conversation escalates, therefore requiring facilitator intervention; and Facilitator Interventions datasets, which include comments by facilitators in active discussions, sometimes annotated with the tactics employed. Some datasets contain information that can be used in multiple tasks. An overview of the surveyed datasets and their categories in our taxonomy can be found in Table 1.
Name Task Size Content Wikipedia Disputes De Kock and Vlachos (2021) Conversation Derailment 7,42574257,4257 , 425 D Includes annotations for several ‘dispute tactics.’ WikiConv Hua et al. (2018) Facilitator Interventions 91,000,0009100000091,000,00091 , 000 , 000 D Includes moderation meta-data such as comment edits and deletions. Conversations Gone Awry Zhang et al. (2018) Conversation Derailment 4,18841884,1884 , 188 D Predicts derailment by analyzing rhetorical tactics, human-annotated. Chang and Danescu-Niculescu-Mizil (2019) (1) Conversation Derailment 4,18841884,1884 , 188 D Extends the ‘Conversations Gone Awry’ dataset. Chang and Danescu-Niculescu-Mizil (2019) (2) Conversation Derailment 6,84268426,8426 , 842 D Based on the r/ChangeMyView subreddit. Park et al. (2012) Conversation Derailment Facilitator Interventions 1,67816781,6781 , 678 C Comprised of 4 datasets. Includes 19 intervention types belonging to 7 moderator roles. RegulationRoom Falk et al. (2021) Conversation Derailment 3,00030003,0003 , 000 C Extends the dataset of Park et al. (2012). DeliData Karadzhov et al. (2021) Facilitator Interventions $500500$ D Group discussions, includes task-oriented quality measure which may be used to approximate discussion quality. Wiki-Tactics De Kock et al. (2022) Facilitator Interventions $213213$ D Based on Wikipedia Disputes, includes moderation action metadata such as comment edits and deletions. UMOD Falk et al. (2024) Facilitator Interventions 2,00020002,0002 , 000 C Based on the r/ChangeMyView subreddit, annotated for facilitation tactics and AQ. Fora Schroeder et al. (2024) Facilitator Interventions $262262$ D Original dataset revolving around experience-sharing, annotated for facilitation tactics.
Table 1: Overview of reviewed datasets. Unnamed datasets are referred to by the names of the authors only. The size reflects the number of annotated conversations, disregarding unlabeled data. D indicates the number of discussions. C indicates the number of individual comments or dialogue turns.
8 LLM Discussion Facilitation Roadmap
Evaluation
LLMs can serve as automated discussion quality annotators (§4). Are these annotators infallible? Not yet. Certain dimensions, especially those that are highly subjective (e.g., pragmatic understanding), remain challenging for LLMs to annotate accurately. But we must take into account that even human annotations tend to be polarized for such subjective quality dimensions Argyle et al. (2023), largely due to sociodemographic background effects and personal biases Beck et al. (2024); Sap et al. (2020).
On the other hand, prompted LLMs offer a more scalable and cost-effective alternative for annotating discussion quality compared to human annotation and traditional (or self-) supervised training on large annotated datasets. Using LLMs for annotation, however, requires careful model selection considering whether models are open or closed source, model size, model alignment, as well as prompt selection, and (if applicable) fine-tuning requirements. These choices should be tailored to the specific quality dimension being evaluated.
Facilitation
Intervention types should be adapted to the different legal frameworks, rules, and social norms of each community/platform. While there are exhaustive surveys on intervention types and policies, such as that of Schaffner et al. (2024), there is yet no methodology to train human or artificial facilitators according to these factors. We posit that experiments using exclusively LLM user/facilitator-agents are necessary to sustainably test new facilitation strategies and interventions per community and platform, as in other NLP tasks that involve LLM-generated conversation Ulmer et al. (2024); Cheng et al. (2024b); Park et al. (2022, 2023), before testing the resulting facilitators in costly experiments with human participants. Finally, the datasets presented in Table 1 can be used to train and assess LLM facilitators in the future, as well as to generate additional data—similar to the existing ones, but with controlled modifications—to stress-test various facilitators in particular settings (e.g., predicting or recovering from a conversation derailment).
9 Conclusions
This survey examined online discussion evaluation and facilitation by bridging insights from Social Science and NLP, with a focus on the growing role of LLMs. We introduced a new discussion evaluation taxonomy, with categories that should remain flexible depending on the evaluation task and the characteristics of the discussion. In terms of intervention strategies, both human- and machine-driven advancements show significant promise in improving the quality of interventions, helping online discussions remain constructive, and resistant to derailment. Most facilitation datasets still originate from human online conversations, with research yet to fully explore the capabilities of LLMs. Taking the above into account, we believe that now is the time to embrace LLMs for facilitation to foster healthier and more constructive conversations.
10 Limitations
This survey is not without its limitations. While we have attempted to present a comprehensive overview of facilitation methods, certain techniques, such as summarization, could be explored in greater depth. Since summarization is a vast subfield of NLP, it was only briefly mentioned.
Moreover, it is important to highlight that most research on facilitation has been conducted solely in English-speaking online spaces. The inherent limitations of LLMs in handling other languages and cultural contexts must be considered. As a result, these findings may not be easily applicable to other regions of the world.
Finally, the majority of real-world online discussions and deliberations happen in the context of communities, where group dynamics (social behaviors, power structures, norms, and interactions) apply. Thus, a fuller review of facilitation would have to account for the internal dynamics of such communities, as well as the wider role of the facilitator as a figure that not only helps in the conversation but has a social status in the group as well.
11 Ethical Considerations
Although AI and LLMs in particular can be effectively used as discussion facilitators, offering dynamic, responsive discussion support, their deployment must meet strict transparency, safety, and accountability standards, especially for high-risk applications, as stated in the EU AI Act. [^1] https://digital-strategy.ec.europa.eu/en/policies/regulatory-framework-ai For example, a person or minority group may have been unfairly disadvantaged in an AI-enhanced deliberation. It is also necessary for the users to be aware that they are interacting with AI facilitators. Ideally the consent of the users should be sought before using any sort of AI-enhanced discussion platform.
Even if LLMs facilitators eventually achieve a high level of autonomy, it is advisable to maintain human oversight. Keeping a human-in-the-loop ensures greater transparency and enables effective error prevention, detection, and correction.
Acknowledgments
This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program.
Appendix A Acronyms
NLP
Natural Language Processing
ML
Machine Learning
LLM
Large Language Model
AM
Argument Mining
ML
Machine Learning
IR
Information Retrieval
AQ
Argument Quality
Appendix B Keywords for Literature Query
Keyword Selection
online discussions, deliberation, dialogue,
discussion evaluation, discussion metrics,
dialogue, deliberation, NLP, AI, discussion quality,
argument mining, survey, LLM, conversation,
moderation, facilitation, communication, democracy
AI dialogue systems, group dynamics
Table 2: Keywords for search engine queries
Appendix C Terminology Background
Here we explain how the surveyed articles were selected. We also explain our reasoning for choosing and disambiguating certain terms (see §2). The definitions of the terms can be found in Table 3.
Facilitation vs. Moderation
“Moderation”, as a term, is more common in Computer Science and NLP, while facilitation is prevalent in Social Sciences Vecchi et al. (2021); Kaner et al. (2007); Trenel (2009). Moderators enforce rules and ensure orderly interactions, usually with the threat of disciplinary action, though they can also act as community leaders Falk et al. (2024); Seering (2020); eRulemaking Initiative (2017). Facilitators, on the other hand, guide discussions, promote participation, and structure dialogue, particularly in online deliberation and education platforms Asterhan and Schwarz (2010). Despite these distinctions, the terms are sometimes used interchangeably Cho et al. (2024); Park et al. (2012); Kim et al. (2021), while it is also common for moderators to use facilitation tactics eRulemaking Initiative (2017); Park et al. (2012); Kim et al. (2021); Cho et al. (2024); Schluger et al. (2022).
Concept Definition and Characteristics
Discussion Broad term encompassing informal and formal exchanges, including online discussions in fora. Can involve elements of debate, dialogue, and deliberation.
Dialogue Collaborative interaction aimed at shared understanding and alignment. Emphasizes cooperation rather than competition. Also refers to dialogue systems in NLP (task-oriented or chatbot conversations).
Deliberation Structured discussion focusing on informed decision-making with reasoned argumentation and diverse perspectives. Less about persuasion, more about collective reasoning.
Debate Adversarial interaction where participants aim to persuade or defend positions rather than achieve mutual understanding. Focused on rhetorical effectiveness.
Thread-style Discussions Online discussions structured in tree/thread formats (e.g., Reddit). Can incorporate elements of all rhetorical styles (debate, dialogue, deliberation).
Discussion Quality Subjective measure influenced by cultural background, engagement, and type of discussion. Defined by socio-dimensional aspects of participant experiences.
Moderation Ensures orderly interactions by enforcing guidelines. Moderators can be volunteers or employees, often associated with disciplinary actions.
Facilitation Encourages equal participation and organizes discussion flow. More common in deliberative and educational contexts, though often used interchangeably with moderation.
Table 3: Definition of terms used in this survey.
Pre-moderation and Post-moderation
Multiple taxonomies have been proposed to describe the temporal dimension of moderation; that is, when moderator action is applied in relation to when the content is visible to the users Veglis (2014); Schluger et al. (2022). These taxonomies are very similar to each other, and usually boil down to the following distinctions:
  • Pre-moderation: The user is dissuaded, or prevented from, posting harmful content. Pre-moderation techniques can include nudges at the writing stage Argyle et al. (2023), reminders about platform rules Schluger et al. (2022), or even a moderation queue where posts have to be approved before being visible to others Schluger et al. (2022).
  • Real-Time: The moderator is part of the discussion and intervenes like a referee would during a match.
  • Ex-post: The moderator is called after a possible incident has been flagged and makes the final call.
Discussion, Deliberation, Dialogue, Debate
There is little to no consensus on how to properly define terms such as “discussion” and “dialogue” Russmann and Lane (2016); Goñi (2024). In this section, we attempt to disambiguate the use of such terms for the purposes of our survey and based on the existing related work. First, our study focuses on discussions, a broader term encompassing various informal and formal exchanges, including online discussions in fora Russmann and Lane (2016), with which we are mainly concerned. In contrast, dialogue refers to collaborative interactions in which participants work toward a shared understanding and alignment Rose-Redwood et al. (2018); Bawden (2021); Goñi (2024). Studies on dialogue emphasize its cooperative nature, aiming for mutual insight rather than competition Bawden (2021). Dialogue can also refer to dialogue systems, a major NLP sub-area, traditionally including both task-oriented dialogues and casual conversation (Eliza-like) [^2] http://web.njit.edu/~ronkowit/eliza.html “chatbots” Liu et al. (2023); Sun et al. (2021).
A more specific concept is deliberation, which involves structured discussions aimed at informed decision-making, often prioritizing reasoned argumentation and the consideration of diverse perspectives Degeling et al. (2015); Lo and McAvoy (2023). Meanwhile, debate is typically adversarial, where participants focus on persuading others or defending their positions. Unlike dialogue or deliberation, debate centers more on winning or convincing, making it less about collective reasoning and more about rhetorical effectiveness Lo and McAvoy (2023). Debates also typically have much stricter (and enforced) rules than other discussions.
For this study, we specifically focus on online written discussions, particularly those occurring in thread- or tree-style formats Seering (2020). A thread is a collection of messages or posts grouped together in an online forum, discussion board, or messaging platform (such as Reddit). It begins with an initial post (often called the original post, or OP), and subsequent replies are ordered either chronologically or by relevance. Threads usually address a specific topic or question and allow users to engage in discussions about that subject. A thread may grow as users contribute more responses. It must be noted, however, that this type of discussion can contain elements from all the other discussion styles. For example, the adversarial element of the debates, or the argumentative element that can be found both in dialogues and deliberations.
Discussion Quality
The success of a discussion is often subjective, influenced by a variety of factors such as the cultural background and linguistic proficiency of the participants Zhang et al. (2018), as well as their level of engagement See et al. (2019). It also depends on the type of the discussion, since some types of discussions, such as deliberations or debates, may not aim at consensus. Given these complexities, we adopt the definition proposed by Raj Prabhu et al. (2021), which views the perceived discussion quality as a measurement that attempts to quantify interactions by taking into account multiple socio-dimensional aspects of individual experiences and abilities.
Appendix D Methodology
The search and article selection of this survey was conducted using specific keywords in academic search engines (e.g., Google Scholar, Semantic Scholar, Scopus), digital libraries and repositories (e.g., ACL Anthology, ACM Digital Library, IEEE Xplore, JSTOR). We focused on peer-reviewed publications written in English between 2014 and 2024, granting exceptions only for established works predating this period. Additionally, we reviewed other cited papers that appeared highly relevant, provided they were peer-reviewed and cited by more than 20 citations of other researchers, unless the topic was very niche, in which case we judged by its content. The search strategy incorporated keywords and phrases related to LLMs, discussion facilitation, and discussion evaluation. The list of keywords used is provided in Table 2. The search was further informed by existing survey articles, such as those by Vecchi et al. (2021) and Wachsmuth et al. (2024), which served as starting points both for identifying relevant literature and for specifying the vocabulary used in the keyword search.
Appendix E Discussion Quality Taxonomy
In this part of the Appendix, we present a table summarizing the discussion evaluation taxonomy (§4). The dimensions are outlined alongside both pre-LLM and LLM-based approaches, while also highlighting their respective contributions to facilitation. The dimensions are color-coded for clarity, with orange indicating associated dimensions that could serve as early signs of potential derailment, green marking signs of constructive growth—i.e., conversations going well or worth participating in—and pink denoting interaction dynamics.
Dimension Facilitation Use Pre-LLM Approaches LLM Approaches Structure & Logic Argument structure &
analysis
Spot claim-evidence chains; raise early-warning flags; keep debate fact-centred Argument-mining pipelines: claim/premise detection; AQ scoring; graph & neural models Zero/few-shot AQ labelling; argument-structure parsing; on-the-fly argument-map summaries
Coherence & flow Detect topic drift; redirect or bridge gaps Entity-grid & sequential coherence models; topic modelling; dialogue state tracking Prompted coherence scoring; chain-of-thought flow checks; off-topic suggestions Turn-taking Monitor balance (entropy/Gini); nudge silent voices; avoid dominance Turn-entropy / Gini metrics; rule-based alarms Context-window turn counts; balanced-participation prompts Language features Track hedges, 2nd-person spikes, jargon; trigger clarification or civility nudges Lexicon features; n-gram-based hedging detectors Style-transfer rephrasers; embedding hedge detection; tone-repair suggestions Speech & dialogue acts Identify interruptions, proposals, question types; score deliberative quality Dialogue-act tagging with ISO/DAMSL labels Few-shot Dialogue Act tagging; tactic selection based on Dialogue Act patterns Pragmatic comprehension Resolve implicatures & sarcasm; surface hidden misunderstandings Commonsense reasoning (Knowledge Base + neural); limited coverage In-context reasoning; auto clarifying questions Social Dynamics Politeness Forecast derailment; issue civility nudges or positive reinforcement Politeness lexicons; domain-independent classifiers Annotation & polite rewrites; policy-violation explanations Power & status Detect dominance; invite low-status voices; rebalance floor Style-matching, pronoun analysis; social-role features Power imbalance estimation; moderator suggestions Disagreement Distinguish constructive vs destructive dissent; de-escalate Graham-hierarchy / stance detection Few-shot labelling; automatic reframing prompts Emotion & Behavior Empathy Encourage empathic turns; highlight emotional cues Lexicon/coding empathy classifiers; affective features Perceived-empathy scoring; supportive paraphrases Toxicity Flag harmful language; decide moderation step BERT/toxicity classifiers; detox lexicons Detection + rewrite suggestions; policy chat Sentiment Track emotional climate; intervene at negativity spikes Lexicon & neural sentiment analysis Prompt-based labelling; tone-shift detection Controversy Sense polarization; invite balancing views Topic-polarity metrics; ideology models Ideology tagging; polarity-aware summaries Constructiveness Stream score; escalate or summarize based on trend Feature-based classifiers (linguistic, discourse) Constructive-rewrite coaching Engagement & Impact Engagement Detect lulls or dominance; prompt interaction Turn/word counts; reply-time gaps Auto-recaps; invite quiet users Persuasion Spotlight evidence-based arguments; dampen manipulation Lexical overlap; ethos/pathos/logos; persuasion prediction Outcome prediction; neutral framing suggestions Diversity & Informativeness Monitor viewpoint spread & info density Topic-diversity indices; IR-based scoring Simulate perspectives; propose links
Table 4: Summary of discussion quality dimensions and corresponding pre-LLM and LLM-based facilitation strategies.
Appendix F Online Discussion Example with Color-coded Politeness Markers
Turn Utterance
0 Why should we help people based on race, and say “we’ll help everyone who’s black, because they could be poor” instead of just “we’ll help everyone who’s poor, in which black people make up a proportionally larger amount”?
1 That study is worse than useless unless it also distinguishes between “black sounding” names that are associated with wealth and poverty.
2 That wouldn’t discount it, that would just add another intersectional axis to investigate. >which I know without looking that it didn’t. How rational.
3 It’s certainly more rational than unquestioningly swallowing everything I read, as some people do. Did this study of yours also test difficult to pronounce Polish names, or Russian names? Or would that have interfered too much with the foregone conclusion they were attempting to reach?
4 Are you implying that’s what I have done? You may be the only one making assumptions here.
Table 5: Dissucssion example from the Reddit Change My View dataset Chang and Danescu-Niculescu-Mizil (2019). Color indicates politeness-related features: hedging, 1st person reference, 2nd person reference, direct questions, negative sentiment and positive sentiment. The annotation was produced with a soon-to-be-released annotation toolkit for discussion evaluation.
This table highlights key politeness-related linguistic features such as hedging, personal references, sentiment, and direct questions. These features are essential in the context of facilitation, where the goal is to guide conversations constructively, maintain safety, and foster mutual understanding. By identifying these elements, the facilitator (human or automatic) can better interpret the tone, intent, and emotional weight of each utterance. For example, detecting hedging or positive sentiment can guide the model to adopt a more collaborative tone, while recognizing negative sentiment or accusatory second-person references may prompt it to de-escalate tension and encourage constructive dialogue.